# UFM: A Simple Path towards Unified Dense Correspondence with Flow

### v1

@@ Appendix C TA-WB Training & Testing Dataset
The geometric sampler generates pairs of rendering directions and source–target cameras based on geometric constraints for viewpoint difference and coarse covisibility check. An overview is presented in Fig. S.1.
@@
After rendering the images, we do additional filtering to ensure their quality. We filter out pairs with any of their images containing more than $10\%$ of over- or under-exposed pixels, and if any of the forward/backward covisibility is less than $20\%$. We further check if the pair is solvable, i.e., does the pair provide enough visual evidence to establish a match? To do this, we warp the target image according to the ground-truth label (similar to Fig. 1, 4) and try to match it to the source image with Superpoint + Lightglue [10, 33, 44]. Since warping is done with ground truth, the matcher should ideally return near-zero pixel displacement in covisible regions. If it does not, the pair lacks enough information to support matching. We retain only pairs with an average matching error below 6 pixels.
@@
We use the geometric sampler to select pairs from the OldScandinavia, Sewerage, Supermarket, DesertGasStation, and PolarSciFi environments in TartanAirV2 [62]. We sample approximately equal numbers of pairs from the angular bins $[0^{\circ},30^{\circ}]$, $[30^{\circ},60^{\circ}]$, and $[60^{\circ},90^{\circ}]$, and allocate roughly half as many pairs to the $[90^{\circ},120^{\circ}]$ bin. Samples of the dataset are provided in Fig. S.2.
@@ Appendix D Training the Refinement
We trained the refinement module on the BlendedMVS, MegaDepth, Habitat, and ScanNet++V2 datasets using image pairs as listed in Table 3.3. We selected these datasets due to their relatively high sub-pixel accuracy. The base model was frozen during this stage, and the refinement module was trained for $30$ epochs with a learning rate of $1\cdot 10^{-4}$. All other optimizer settings are the same as the 560 base model training, as detailed in Section 3.4. Fig. S.4 shows a visualization of the trained features, where we see high-frequency and edge-following behavior that encodes the local details.
@@ Appendix E Effect of Freezing the Encoder
We found that freezing the DINOv2 encoder and using its last-layer features was suboptimal for UFM. Specifically, when training UFM on the FlyingChairs dataset, we observed a significant validation EPE gap between using features from the last layer versus intermediate layers of the frozen DINOv2 encoder. As shown in Fig. S.5, UFM trained with the last layer features from frozen DINOv2 obtained near $3$ EPE, whereas features from layer 10 yielded sub-pixel performance. This gap is not observed in the finetuned setting, given sufficient training.
@@ E.2.  Probing Experiment:
The outline of our probing experiment is shown in Fig. S.6. We select a relatively small dataset and disable all augmentation to ensure that training will converge. We infer each pair of images through the frozen DINOv2 encoder and project the source and target features through a layer-specific linear layer. We then compute softmax dot-product similarity to mimic the global attention mechanism. Each layer’s probe is trained independently, and its performance reflects how well the layer encodes corresponding features that can be revealed during the global attention. Patch-wise similarity is defined as the proportion of pixel-wise correspondences between patches, weighted by covisibility. Formally, given correspondence and covisibility labels $\phi^{gt},C^{gt}$, the ground-truth patch similarity $s(P_{s},P_{t})$ between a source patch $P_{s}\subset I_{1}$ and target patch $P_{t}\subset I_{2}$ is defined as:
@@ E.3.  Probing Results using Correlation:
To test whether the cross-entropy loss from probing correlates with EPE performance, we trained UFM using different frozen layers of DINOv2 on the FlyingChairs dataset and collected their loss in the probing experiment. We normalized the cross-entropy loss value into probing performance between $[0,1]$. According to Fig. S.7, we found a strong correlation between the loss in the probing experiment and the final validation EPE, and the peaks differ only by $2$ of the total $24$ layers. This suggests that, for UFM, probing performance may serve as a reliable indicator for selecting effective feature layers. Furthermore, this supports the hypothesis that the last layer of DINO does not provide the strongest corresponding feature, thus leading to suboptimal performance. We further show additional probing results on other datasets, resolutions, and DINOv2 encoder sizes in Fig. S.8. We found a consistent trend where the intermediate layers encode stronger correlating features and perform better.

### v2

@@ Appendix C TA-WB Training & Testing Dataset
The geometric sampler generates pairs of rendering directions and source–target cameras based on geometric constraints for viewpoint difference and coarse covisibility check. An overview is presented in Fig.˜S.1.
@@
After rendering the images, we do additional filtering to ensure their quality. We filter out pairs with any of their images containing more than $10\%$ of over- or under-exposed pixels, and if any of the forward/backward covisibility is less than $20\%$. We further check if the pair is solvable, i.e., does the pair provide enough visual evidence to establish a match? To do this, we warp the target image according to the ground-truth label (similar to Fig.˜1, 4) and try to match it to the source image with Superpoint + Lightglue [10, 38, 49]. Since warping is done with ground truth, the matcher should ideally return near-zero pixel displacement in covisible regions. If it does not, the pair lacks enough information to support matching. We retain only pairs with an average matching error below 6 pixels.
@@
We use the geometric sampler to select pairs from the OldScandinavia, Sewerage, Supermarket, DesertGasStation, and PolarSciFi environments in TartanAirV2 [68]. We sample approximately equal numbers of pairs from the angular bins $0^{\circ},30^{\circ}]$, $30^{\circ},60^{\circ}]$, and $60^{\circ},90^{\circ}]$, and allocate roughly half as many pairs to the $90^{\circ},120^{\circ}]$ bin. Samples of the dataset are provided in Fig.˜S.2.
@@ Appendix D Training the Refinement
We trained the refinement module on the BlendedMVS, MegaDepth, Habitat, and ScanNet++V2 datasets using image pairs as listed in Table 1. We selected these datasets due to their relatively high sub-pixel accuracy. The base model was frozen during this stage, and the refinement module was trained for $30$ epochs with a learning rate of $1\cdot 10^{-4}$. All other optimizer settings are the same as the 560 base model training, as detailed in Section 3.4. Fig. S.4 shows a visualization of the trained features, where we see high-frequency and edge-following behavior that encodes the local details.
@@ Appendix E Effect of Freezing the Encoder
We found that freezing the DINOv2 encoder and using its last-layer features was suboptimal for UFM. Specifically, when training UFM on the FlyingChairs dataset, we observed a significant validation EPE gap between using features from the last layer versus intermediate layers of the frozen DINOv2 encoder. As shown in Fig.˜S.5, UFM trained with the last layer features from frozen DINOv2 obtained near $3$ EPE, whereas features from layer 10 yielded sub-pixel performance. This gap is not observed in the finetuned setting, given sufficient training.
@@ E.2. Probing Experiment:
The outline of our probing experiment is shown in Fig.˜S.6. We select a relatively small dataset and disable all augmentation to ensure that training will converge. We infer each pair of images through the frozen DINOv2 encoder and project the source and target features through a layer-specific linear layer. We then compute softmax dot-product similarity to mimic the global attention mechanism. Each layer’s probe is trained independently, and its performance reflects how well the layer encodes corresponding features that can be revealed during the global attention. Patch-wise similarity is defined as the proportion of pixel-wise correspondences between patches, weighted by covisibility. Formally, given correspondence and covisibility labels $\phi^{gt},C^{gt}$, the ground-truth patch similarity $s(P_{s},P_{t})$ between a source patch $P_{s}\subset I_{1}$ and target patch $P_{t}\subset I_{2}$ is defined as:
@@ E.3. Probing Results using Correlation:
To test whether the cross-entropy loss from probing correlates with EPE performance, we trained UFM using different frozen layers of DINOv2 on the FlyingChairs dataset and collected their loss in the probing experiment. We normalized the cross-entropy loss value into probing performance between $[0,1]$. According to Fig.˜S.7, we found a strong correlation between the loss in the probing experiment and the final validation EPE, and the peaks differ only by $2$ of the total $24$ layers. This suggests that, for UFM, probing performance may serve as a reliable indicator for selecting effective feature layers. Furthermore, this supports the hypothesis that the last layer of DINO does not provide the strongest corresponding feature, thus leading to suboptimal performance. We further show additional probing results on other datasets, resolutions, and DINOv2 encoder sizes in Fig.˜S.8. We found a consistent trend where the intermediate layers encode stronger correlating features and perform better.
@@
Appendix F Discussion on the use of Matching in UFM Architecture
As many prior works have shown the benefits of using matching for correspondence prediction, we would like to further discuss and ablate our choice to regress the initial flow combined with applying matching only at pixel resolution, not patch resolution.
Matching at the patch level
Matching in patch-level aggregates subsets of pixels as a patch, then matches them with other patches to produce an initial flow hypothesis at low resolution. The initial coarse flow will be upsampled and refined later, following the “coarse-to-fine” paradigm. To better present our findings, we would first like to clarify two possible meanings of the term “coarse-to-fine”: (1) the predicted flow is upsampled from a coarse to a fine spatial resolution, or (2) a rough prediction is progressively refined for higher accuracy. While prior methods often combine both, our approach uses (2) but not (1).
To support our choice, we replaced the DPT head of UFM 560 with the UniMatch matching head and trained with the same protocol as described in Section˜3.4. We adopt this head as a representative of patch-scale matching methods. Hence, the matching head now benefits from the same backbone capacity and unified dataset as UFM 560. The only differences are the final flow prediction head and how the flow loss is applied. To maintain equivalence in supervision, we modified the UFM robust loss to match how UniMatch applies its loss - supervising on both the final flow and the bilinearly upsampled coarse flow. Essentially, the UniMatch head first performs patch-level (coarse) matching, which is then convex upsampled to pixel-wise correspondence. We follow their original codebase and refer to formulas (1) - (7) from GMFlow [73] for the exact definition of the UniMatch head.
We evaluated them on covisible end-point-errors and report the results in Table˜S.2, where we observed that the UniMatch Head yields a $22\%$ increase in EPE on average. Hence, we conclude that the regression head is more accurate in predicting correspondence and scales better to the combined dataset, corroborated by evidence in Fig.˜5.
Table S.2: Matching Head Ablation: With identical backbone and training data, a regression-based flow head (UFM design) using DPT achieved higher precision than a patch-wise matching head from UniMatch.
Head Type Sintel-C Sintel-F KITTI DTU ETH3D TA-WB
EPE [covis]
Patch Matching + Upsample 0.91 1.49 2.05 5.36 5.47 13.85
Regression (ours) 0.79 1.44 1.87 2.64 5.56 12.87
Matching at the pixel level
Direct global matching of pixel descriptors allows MASt3R to achieve high precision even for large motions. Using a similar architecture and dataset scale, we show that the regression paradigm in UFM 560 achieves competitive EPE performance on challenging wide-baseline datasets, where the average flow reaches about 90 - 200 pixels. We attribute this to UFM’s superior supervision efficiency compared to MASt3R. While MASt3R trains with at most 4096 correspondence pairs per frame, UFM leverages supervision from all covisible pixels. This dense supervision is a key factor that allows UFM to achieve similar or better accuracy with much higher speed than MASt3R.
Matching and Regression are not mutually exclusive
Through UFM-Refine, we demonstrate that matching-based local refinement can be effectively combined with a regressed initial hypothesis to further enhance accuracy. This forms a key distinction between UFM-Refine and previous dense correspondence approaches. Only with an efficient and accurate regressed initialization can one confidently narrow the search range for pixel-wise matching, thereby achieving both the efficiency of regression and the precision of pixel-wise matching.
Appendix G Additional Evaluations
We provide additional zero-shot evaluations of UFM on Sintel and KITTI test sets in Table S.3, S.4. While we still observe UFM outperforms UniMatch under zero-shot settings (neither method was trained on Sintel or KITTI), we observed UFM is biased towards static objects. As in Table˜S.4, UFM is significantly better in the background than in the foreground objects, which are mostly dynamic. Future work may include more sceneflow datasets, such as ParallelDomain [63], to balance the distribution.
Table S.3: Zero-Shot Performance on Sintel Test Set: UFM outperforms UniMatch in the overall metric and most sub-items, except for sub-items on Sintrel-Clean.
Dataset Method EPE all EPE matched EPE unmatched d0-10 d10-60 d60-140 s0-10 s10-40 s40+
UFM 980 - refine 3.28 1.75 15.80 3.28 1.27 1.21 0.69 2.07 19.07
Sintel-Final UniMatch-CT 4.14 1.97 21.86 3.77 1.59 1.28 0.89 2.43 24.40
UFM 980 - refine 1.68 0.72 9.52 1.54 0.67 0.43 0.48 1.23 8.71
Sintel-Clean UniMatch-CT 1.80 0.57 11.85 1.18 0.45 0.40 0.37 0.93 11.03
Table S.4: Zero-Shot Performance on KITTI Test Set: UFM outperforms UniMatch in the overall metric and background objects, but not foregrond objects, which are mostly dynamic.
Method Range F1-bg F1-fg F1-all
UFM 980 - refine All 8.67 22.08 10.9
UniMatch-CT All 18.01 17.27 17.89
UFM 980 - refine Not occluded 4.93 19.6 7.59
UniMatch-CT Not occluded 9.75 14.61 10.63
Appendix H Visualization on the Robust Charbonnier Loss
We provide an additional visualization of the robust Charbonnier loss. The loss is defined as:
$L_{robust}\left(x;\alpha,c\right)=\frac{\left|\alpha-2\right|}{\alpha}\left(\left(\frac{\left(\frac{x}{c}\right)^{2}}{\left|\alpha-2\right|}+1\right)^{\frac{\alpha}{2}}-1\right)$ (S.9)
We followed RoMa’s choice of $\alpha=0.5$, $c=0.24$. As shown in  Fig.˜S.9, this combination of $\alpha$ and $c$ puts a high gradient on inliers while maintaining non-negligible gradient for even the largest outliers to slowly correct them.
Figure S.9: Gradient of the Robust Charbonnier Loss: Compared to $L_{1}$ loss, the robust loss used by UFM puts higher gradient to the inliers (peaks at error of $0.42$ pixel), while maintaining non-negligible ($0.12$ at $500$ pixel error) gradient for even the largest outliers to slowly correct them.